OpenStudy (yb1996):

Use the definition of a derivative to find dy/dx of y=f(x)^2.

OpenStudy (yb1996):

I know that by taking the derivative the answer will be y' = 2f(x)f'(x), but how do I get that by using the definition of a derivative?

OpenStudy (phi):

by the definition, do you mean limit f(x+h)- f(x) / h ?

OpenStudy (yb1996):

Yes

OpenStudy (mathmale):

Just verifying what y ou have there: Do you want to find the derivative, with respect to x, of y = [f(x)]^2 OR y=x^2?

OpenStudy (yb1996):

y = [f(x)]^2

OpenStudy (mathmale):

Great. Be sure to use the Chain Rule as well as the Power Rule. By the way, you're in great hands with phi.

OpenStudy (phi):

I think you are asking to use \[ \lim_{h\rightarrow 0} \frac{ g(x+h) - g(x)}{h} \] where g(x) = \(f^2(x) \)

OpenStudy (yb1996):

Yes

OpenStudy (phi):

we get \[ \frac{f^2(x+h)-f^2(x)}{h} \\ \frac{(f(x+h)-f(x))\cdot (f(x+h)+f(x)}{h} \] is that enough of a hint ?

OpenStudy (phi):

remember the limit of a product is the product of limits so you can write that as \[ \lim_{h\rightarrow 0} \frac{f(x+h)-f(x)}{h} \cdot \lim_{h\rightarrow 0} (f(x+h)+f(x)) \]

OpenStudy (phi):

for the sum part, the limit of a sum is the sum of the limits

OpenStudy (yb1996):

Ok, so that would simplify to f'(x) * 2f(x)

OpenStudy (phi):

yes

OpenStudy (yb1996):

Ok, thank you for you help!

OpenStudy (phi):

Though I am thinking there must be another way to do this, because that trick does not work for e.g. f^3(x) (or at least I don't see how)

OpenStudy (yb1996):

Probably